OpenStudy (anonymous):

Probability and Statistics: Gaussian Distribution: Suppose that a certain sample space of numbers can be modeled by a Gaussian distribution with a mean of 68 and a standard deviation of 2.5. First, find the percentage of the sample space whose numbers is between 66 and 71. Second, find the percentage whose numbers is approximately 72. -- For the first one, I tried to normalize 68 into -0.8 and 71 to 1.2 (using (parameter-mean)/(std. dev)), then I tried to numerically integrate the standard Gaussian PDF from -0.8 to 1.2 to get 67.31%. Is this right?

OpenStudy (anonymous):

I'm not sure how to do the second one, though..

OpenStudy (kropot72):

I get 67.30% for the first one using a standard normal distrubution table. I will post a method and a solution for the second one. Please wait a few minutes.

OpenStudy (anonymous):

Thank you very much.

OpenStudy (kropot72):

The percentage of numbers approximately equal to 72 can be found by appropriate calculation of the z-scores for X above and below 72. For example if the numbers lie within 72 plus and minus 0.5: \[z _{1}=\frac{71.5-68}{2.5}\] \[z _{2}=\frac{72.5-68}{2.5}\] Using these values of z-score the required percentage = 4.49%

OpenStudy (anonymous):

How did you decide to use 0.5 though? Is this number arbitrary?

OpenStudy (kropot72):

Good question! 0.5 is a correction used when changing from a discrete to a continuous distribution such as in the normal approximation of the binomial distribution. Although it is not directly relevant to your question, I thought it would be a good starting point. You can try tightening the limits by using smaller values than 0.5 and assessing the senstitivity if you like.

OpenStudy (anonymous):

This is really late, but thank you very much for your help!

OpenStudy (kropot72):

You're welcome :)